# The DECADE cosmic shear project III: validation of analysis pipeline using spatially inhomogeneous data

### v1

@@ 6.3 Object property splits & results
We now repeat the analysis of Section 6.2 but splitting the sample on galaxy properties rather than through a survey mask. While this is not a strict test of spatial inhomogeneities, the property splits will isolate subsets of the catalog that may be more susceptible to inhomgeneity in various observing conditions; for example, the shapes and photometry of galaxies with small sizes can be more strongly affected by PSF variations. All properties used in the splits are taken from the MetaCalibration catalog, as this allows any selection-based bias in the shear to be robustly calibrated through the MetaCalibration formalism (e.g., Sheldon & Huff 2017). We split on colors, magnitudes, signal-to-noise ratios, and sizes. The color split is motivated by Samuroff et al. (2019) and McCullough et al. (2024) who split the DES Y1 and Y3 sample, respectively, into red and blue galaxies and redo the cosmic shear analysis for each split. Their color split (see Table 1 in McCullough et al. (2024)) is a $r-z$ cut that depends on the tomographic bin. Given the setup of our pipeline, we only perform a non-tomographic $r-z$ split across the full catalog. We choose $r-z\gtrless 0.85$ as it is the median value, across the four bins, used in McCullough et al. (2024). We then consider a range of magnitude cuts and size cuts, which probe depth-based and PSF-based effects, respectively. The signal-to-noise cut is correlated with both magnitude and size-based cuts. All the non-color selections were defined using the median value of their distributions in the fiducial shape catalog and the chosen values are listed alongside our results.
Figure 6 shows the constraints from the different splits. Similar to the test in Section 6, most (all) constraints from the subsets are within $1\sigma$ ($2\sigma$) of the “Fiducial” estimate. Each subset and its complement, for all splits, are also consistent within $2\sigma$ of each other. These results showcase the robustness of our cosmology results against systematics that preferentially affect certain types of objects in the catalog (e.g., small objects) and further highlight that the cosmic shear analysis pipeline can handle data of a less pristine nature.
@@ 7 Summary
In this paper, we outline the methodology used to perform the cosmological inference of the Dark Energy Camera All Data Everywhere (DECADE) shear catalog. The modeling framework and implementation builds heavily on the DES Year 3 cosmic shear analysis (Secco & Samuroff et al. 2022; Amon et al. 2022), and incorporates recent updates as well (DES and KiDS Collaborations et al. 2023). After describing the basic modeling choices, we perform a series of specific tests to: (1) ensure that our baseline analysis choices are robust to different choices in the model, and; (2) stress-test our pipeline by performing the same analysis on forty-six different subsets of the dataset, each defined as a selection on observing conditions or galaxy properties. Some of these tests are particularly important as the DECADE dataset is a compilation of very different DECam observing campaigns and thus, the characteristics of this dataset are more inhomogeneous than conventional weak lensing surveys.
@@
  • For the covariance matrix, we use CosmoCov to generate an analytical covariance matrix that includes both Gaussian and non-Gaussian contributions. We also specifically test the shape noise model for the covariance matrix using simulations and find that the inhomogeneity in the galaxy distribution (at the level seen in DECADE) does not appreciably impact the model predictions.
  • Our expected constraint power is similar to the DES Y3 “Fiducial” analysis constraints but 50% worse than the DES Y3 “Optimal” analysis constraints, given the latter removes significantly fewer scales in its analysis.
@@
  • This is the first time the above end-to-end test is done with such completeness in a weak lensing survey. This was made possible by self-consistently recalibrating the shear estimates, remeasuring the redshift distributions and calibrations, re-deriving the covariance matrix estimate, and re-measuring the data vectors for each subset of the data. We find these tests cover a wide range of systematics in the catalog and recommend future surveys setup infrastructure to enable easy repetition of the processes in their end-to-end pipeline.
This work is part of a series of papers. Paper I details the shear catalog and all associated systematic checks while Paper II describes the redshift distribution and its associated validations/calibrations. The cosmological constraints from the DECADE data, derived using the full analysis methodology discussed here, are presented in Paper IV.
Finally, we re-iterate the last point from our summary: there is significant value in building analysis frameworks that can easily ingest object selection functions defined by a user, and then use it in all steps of the end-to-end cosmic shear pipeline. In this work, we have discussed such functionality in the context of systematics validation, where it is a vital tool in checking that different subsets of the catalog produce statistically consistent cosmology constraints. However, we note that such functionality inherently increases the power and applicability of these datasets to the broader community. For example, combinations of lensing data with other datasets may benefit from using different galaxy selection functions, or different definitions for the tomographic bins. For example, additional selections on the source galaxies can improve the lensing-based mass estimate of galaxy clusters (e.g., Rau et al. 2024), while cross-correlations with other surveys/datasets often use only a subset of the photometric survey (e.g., Shin et al. 2019; Omori et al. 2023; Chang et al. 2023; Sánchez et al. 2023; Anbajagane et al. 2024b), and this can be done more robustly by re-calibrating the relevant subset of the data. Even in lensing-only analyses, we have found uses for alternative tomographic binning of the data (Secco et al. 2022), and also for using different selection functions to isolate data that may be uniquely robust to certain systematics (Samuroff et al. 2019; McCullough et al. 2024). Building such functionality into our future pipelines will significantly improve the usability of lensing data, and enable the broader astrophysics and cosmology community to take better advantage of our precision datasets.
@@ Acknowledgements
All analysis in this work was enabled greatly by the following software: Pandas (McKinney 2011), NumPy (Van der Walt et al. 2011), SciPy (Virtanen et al. 2020), and Matplotlib (Hunter 2007). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.
@@
All catalogs and derived data products (data vectors, redshift distributions, calibrations etc.) will be made public upon acceptance of the cosmology results.
@@ Appendix A Impact of PSF contamination and B-modes
1 $0.0025\pm 0.0070$ $1.6347\pm 0.1630$ $-0.8548\pm 2.2794$
2 $-0.0136\pm 0.0077$ $1.7490\pm 0.1890$ $-3.2707\pm 2.0824$
3 $-0.0058\pm 0.0100$ $1.6712\pm 0.1818$ $-2.8773\pm 2.9182$
4 $-0.0015\pm 0.0107$ $1.8128\pm 0.2270$ $0.8296\pm 3.0527$
Table 3: The PSF coefficients per tomographic bin, as used in Equation (A.1), computed using galaxies from only a given tomographic redshift bin. All measured coefficients are statistically consistent with expected values of $\alpha\lesssim 10^{-3}$ and $\beta,\eta\sim 1$. The method for computing the coefficients follow that describe in Paper I (see their Section 4.5).
Figure 7: The impact of PSF contamination on the $S_{8}$-$\Omega_{\rm m}$ contours. We convert the distribution of $\alpha$, $\beta$, $\eta$ parameters (Equation A.1) into additive offsets to the data vector, $\delta\xi_{\pm}^{\rm\,psf}$, and contaminate a simulated data vector with the ±$\pm 3\sigma$ values of the $\delta\xi_{\pm}^{\rm\,psf}$, distribution. The cosmology constraints are robust even after assuming extreme values of the PSF contamination.
Figure 8: The angular power spectra of $B$-modes in the shape catalog, measured using galaxies from different tomographic bins (see legend). The measurements for all tomographic bin combinations are consistent with no $B$-mode signal. For brevity, we show only a subset of the measurements with the bin-combinations denoted in the legend.
Figure 10: The mean redshift distributions for the four tomographic bins (columns) for some of the splits shown in Section 6.2 and 6.3. The fiducial redshift distribution is shown in gray. The text in the top shows the mean redshift, as well as the uncertainty on it as calibrated by the pipeline in Paper II. The text also denotes the split corresponding to each color. The area-based splits only lead to minimal changes in the mean $n(z)$ — though they may have greater effects in causing spatially varying $n(z)$, which will not be highlighted in this figure which presents the area-averaged $n(z)$ — while the object-based splits cause a clear separation in redshift.
@@ A.1 PSF contamination
We denote the measured two-component galaxy ellipticity in bin $a$, as $\boldsymbol{e}^{\rm obs}_{a}$. This ellipticity will have some contribution from untreated PSF contamination, $\boldsymbol{e}^{\rm obs}_{a}=\boldsymbol{e}^{\rm true}_{a}+\delta\boldsymbol{e}^{\rm psf}_{a}$. Now we can write the two-point correlation of the shapes as,
@@
where we follow previous shape measurement literature in assuming $\boldsymbol{e}^{\rm true}$ does not correlate with $\boldsymbol{e}^{\rm psf}$, i.e., we assume $\langle\boldsymbol{e}^{\rm true}\boldsymbol{e}^{\rm psf}\rangle=0$.
@@
for the same reasons mentioned above, $\boldsymbol{e}^{\rm true}$ does not correlate with $\boldsymbol{e}^{\rm psf}$ so the second and third term average to 00 given $\langle\boldsymbol{e}^{\rm true}\rangle=0$.
@@
where the coefficients $\alpha_{a},\beta_{a},\eta_{a}$ are measured using only the galaxies in tomographic bin $a$. The quantities $\boldsymbol{e}^{\rm model},\boldsymbol{e}^{\star}$, etc. are not dependent on tomographic bin, but the coefficients associated with a given term does depend on the bin. We can write the cross-term, $\langle\delta\boldsymbol{e}^{\rm psf}_{a}\delta\boldsymbol{e}^{\rm psf}_{b}\rangle$ as,
@@
We compute these bin-dependent coefficients using the same techniques as Paper I (see their Section 4.5) but now with only the source galaxies in tomographic bin $a$. The coefficient estimates are listed in Table 3, and are consistent with the non-tomographic case presented in Paper I. We then compute the PSF contamination to the shear two-point data vector, $\delta\xi_{\pm}^{\rm\,psf}$, using Equation (A.1) and the coefficients of Table 3. For each bin, we generate a distribution of $\delta\xi_{\pm}^{\rm\,psf}$ contributions. This is done by sampling from the joint posterior of the $\alpha,\beta,\eta$ parameters in that bin, and generate a $\delta\xi_{\pm}^{\rm\,psf}$ per sample. We then compute the $99.7\%$ upper/lower bounds on $\delta\xi_{\pm}^{\rm\,psf}$ and use those two estimates to contaminate the simulated data vector. The constraints from the contaminated data vectors are shown in Figure 7. The shifts in the cosmology are completely negligible, indicating that PSF contamination has negligible impact on the DECADE cosmology constraints.
@@ A.2 B-modes
In Paper I, we verify that the DECADE shape catalog does not contain any $B$-mode signal to within the statistical precision of the data. We now repeat this exercise, but perform the measurement for different tomographic bin combinations. Here, we have used the Namaster (Alonso et al. 2019) package to measure the B-modes power spectra in our data.
@@
Figure 8 shows the results. We find no evidence of B-modes in any tomographic bin combination. Given we have no detection of a B-mode contamination, we do not estimate a $\delta\xi_{\pm}^{\,\rm B-mode}$ contribution and propogate the impact of a $B$-mode contaminant to cosmology. This follows the choices in Amon et al. (2022). Note that the harmonic-space estimator here differs from the real-space estimator used in Paper I. The former was chosen for this test as our original purpose was to contaminate the simulated data vector with $B$-modes (assuming a $B$-mode signal was detected) and this task requires as input the angular power spectra of $B$-modes. We have also verified that the non-tomographic measurement of the $B$-modes from Namaster is statistically consistent with no signal, which reproduces the result from the real-space estimator in Paper I.
Appendix B Addition results from Survey Property splits

### v2

@@ 6.3 Object property splits & results
We now repeat the analysis of Section 6.2 but splitting the sample on galaxy properties rather than through a survey mask. While this is not a strict test of spatial inhomogeneities, the property splits will isolate subsets of the catalog that may be more susceptible to inhomogeneity in various observing conditions; for example, the shapes and photometry of galaxies with small sizes can be more strongly affected by PSF variations. All properties used in the splits are taken from the MetaCalibration catalog of Paper I, as this allows any selection-based bias in the shear to be robustly calibrated through the MetaCalibration formalism (e.g., Sheldon & Huff 2017). We split on colors, magnitudes, signal-to-noise ratios, and sizes. The color split is motivated by Samuroff et al. (2019) and McCullough et al. (2024) who split the DES Y1 and Y3 sample, respectively, into red and blue galaxies and redo the cosmic shear analysis for each split. Their color split (see Table 1 in McCullough et al. (2024)) is a rzr-z cut that depends on the tomographic bin. Given the setup of our pipeline, we only perform a non-tomographic rzr-z split across the full catalog. We choose $r-z\gtrless 0.85$ as it is the median value, across the four bins, used in McCullough et al. (2024). We then consider a range of magnitude cuts and size cuts, which probe depth-based and PSF-based effects, respectively. The signal-to-noise cut is correlated with both magnitude and size-based cuts. All the non-color selections were defined using the median value of their distributions in the fiducial shape catalog and the chosen values are listed alongside our results.
Figure 6 shows the constraints from the different splits. Similar to the test in Section 6.2, most (all) constraints from the subsets are within $1\sigma$ ($2\sigma$) of the “Fiducial” estimate. Each subset and its complement, for all splits, are also consistent within $2\sigma$ of each other. These results showcase the robustness of our cosmology results against systematics that preferentially affect certain types of objects in the catalog (e.g., objects with small angular sizes) and further highlight that the cosmic shear analysis pipeline can handle data of a less pristine nature.
@@
These split tests are a powerful method for inferring the impact of systematics on cosmology through a completely data-driven approach. Throughout this analysis we did not have to produce any simulations assuming a given systematic effect (and therefore did not need an a-priori understanding of how a given systematic effect impacts the data). The disadvantage of this method is its relative insensitivity — it is best at indicating whether a systematic is causing a $3\sigma$ parameter shift. While this is still valuable for catching hidden or poorly understood systematics, it is not ideal for characterizing systematics that have smaller effects on the posterior constraints. One obvious extension to our approach is to link together the parameter shifts from different splits and determine consistency in cosmology constraints in an ensemble sense, i.e. by summarizing the distribution of shifts (across different splits) through a $p$-value metric. This requires a precise, robust covariance matrix for the correlation function measurements of the different subsets. One approach to do so is using simulation-based estimates. We have not pursued this given the computational cost. However, the calculation is tractable and so could be pursued in the future.
@@ 7 Summary
In this paper, we outline the methodology used to perform the cosmological inference of the Dark Energy Camera All Data Everywhere (DECADE) shear catalog. Our modeling framework and implementation builds heavily on the DES Year 3 cosmic shear analysis (Secco & Samuroff et al. 2022; Amon et al. 2022), and also incorporates recent updates (DES and KiDS Collaborations et al. 2023). After describing the basic modeling choices, we perform a series of specific tests to: (1) ensure that our baseline analysis choices are robust to different choices in the model, and; (2) stress-test our pipeline by applying the same analysis pipeline on forty-six different subsets of the dataset, each defined as a selection on observing conditions or galaxy properties. Some of these tests are particularly important as the DECADE dataset is a compilation of very different DECam observing campaigns and thus, the characteristics of this dataset are more inhomogeneous than conventional weak lensing surveys.
@@
  • For the covariance matrix, we use CosmoCov to generate an analytical covariance matrix that includes both Gaussian and non-Gaussian contributions. We also specifically test the shape noise model for the covariance matrix using simulations, in Appendix C, and find that the inhomogeneity in the galaxy distribution (at the level seen in DECADE) does not appreciably impact the model predictions.
  • Our expected constraint power is similar to the DES Y3 “Fiducial” analysis constraints but 30% worse than the DES Y3 “Optimal” analysis constraints, given the latter removes significantly fewer scales in its analysis (Section 5).
@@
  • This is the first time the above end-to-end test is done with such completeness in a weak lensing survey. This was made possible by self-consistently recalibrating the shear estimates, remeasuring the redshift distributions and calibrations, rederiving the covariance matrix estimate, and remeasuring the data vectors for each subset of the data. We find these tests cover a wide range of systematics in the catalog and recommend future surveys setup infrastructure to enable easy repetition of the processes in their end-to-end pipeline.
This work is part of a series of papers. Paper I details the shear catalog and all associated systematic checks while Paper II describes the redshift distribution and its associated validations/calibrations. The cosmological constraints from the DECADE data, derived using the full analysis methodology discussed here, are presented in Paper IV. Results from combining this DECADE dataset with DES Y3, and with another $\approx 3,\!300\deg^{2}$ of DECADE data in the southern Galactic cap, can be found in Anbajagane & Chang et al. (2025d).
Finally, we reiterate the last point from our summary: there is significant value in building analysis frameworks that can easily ingest object selection functions defined by a user, and then use it in all steps of the end-to-end cosmic shear pipeline. In this work, we have discussed such functionality in the context of systematics validation, where it is a vital tool in checking that different subsets of the catalog produce statistically consistent cosmology constraints. However, we note that such functionality inherently increases the power and applicability of these datasets to the broader community. For example, combinations of lensing data with other datasets may benefit from using different galaxy selection functions, or different definitions for the tomographic bins. Additional selections on the source galaxies can also improve the lensing-based mass estimate of galaxy clusters (e.g., Rau et al. 2024), while cross-correlations with other surveys/datasets often use only a subset of the photometric survey (e.g., Shin et al. 2019; Omori et al. 2023; Chang et al. 2023; Sánchez et al. 2023; Anbajagane et al. 2024b), and this can be done more robustly by recalibrating the relevant subset of the data. Even in lensing-only analyses, we have found uses for alternative tomographic binning of the data (Secco et al. 2022; Campos et al. 2023), and also for using different selection functions to isolate data that may be uniquely robust to certain systematics (Samuroff et al. 2019; McCullough et al. 2024). Building such functionality into our future pipelines will significantly improve the usability of lensing data, and enable the broader astrophysics and cosmology community to take better advantage of our precision datasets.
@@ Acknowledgements
This research uses services or data provided by the Astro Data Lab, which is part of the Community Science and Data Center (CSDC) Program of NSF NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA), Inc. under a cooperative agreement with the U.S. National Science Foundation.
@@
All analysis in this work was enabled greatly by the following software: Pandas (McKinney 2011), NumPy (Van der Walt et al. 2011), SciPy (Virtanen et al. 2020), Matplotlib (Hunter 2007), and Getdist (Lewis 2019). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.
@@
All catalogs and derived data products (data vectors, redshift distributions, calibrations etc.) for the cosmology analysis are now publicly available through the Noirlab Datalab portal (Fitzpatrick et al. 2014; Nikutta et al. 2020) as well as through Globus and other avenues. Please visit dhayaaanbajagane.github.io/data_release/decade for a list of the available dataproducts and their corresponding data access. Our intention is to make all useful products immediately available to the community. Please reach out to DA if a data product of interest to you is not on the above list.
@@ Appendix A Impact of PSF contamination and $B$-modes
1 $0.0025\pm 0.0070$ $1.6347\pm 0.1630$ $0.8548\pm 2.2794$
2 $0.0136\pm 0.0077$ $1.7490\pm 0.1890$ $3.2707\pm 2.0824$
3 $0.0058\pm 0.0100$ $1.6712\pm 0.1818$ $2.8773\pm 2.9182$
4 $0.0015\pm 0.0107$ $1.8128\pm 0.2270$ $0.8296\pm 3.0527$
Table 3: The PSF coefficients per tomographic bin, as used in Equation (A.1), computed separately for each tomographic redshift bin using only galaxies in that bin. All measured coefficients are statistically consistent with expected values of $\alpha\lesssim 10^{-3}$ and $\beta,\eta\sim 1$. The method for computing the coefficients follow that describe in Paper I (see their Section 4.5).
Figure 7: The impact of PSF contamination on the $S_{8}$-$\Omega_{\rm m}$ contours. We convert the distribution of $\alpha$, $\beta$, $\eta$ parameters (Equation A.1) into additive offsets to the data vector (DV), $\delta\xi_{\pm}^{\rm\,psf}$, and contaminate a simulated data vector with the $\pm 3\sigma$ values of the $\delta\xi_{\pm}^{\rm\,psf}$ distribution. The cosmology constraints are robust even after assuming extreme values of the PSF contamination.
Figure 8: The angular power spectra of $B$-modes, $C_{\ell}^{BB}$, in the shape catalog, measured using galaxies from different tomographic bins (see legend). The measurements for all tomographic bin combinations are consistent with no $B$-mode signal. For brevity, we show only a subset of the measurements with the bin-combinations denoted in the legend.
@@ A.1 PSF contamination
We denote the measured two-component galaxy ellipticity in bin $a$, as $\boldsymbol{e}^{\rm obs}_{a}$. This ellipticity will have some contribution from untreated PSF contamination, $\boldsymbol{e}^{\rm obs}_{a}=\boldsymbol{e}^{\rm true}_{a}+\delta\boldsymbol{e}^{\rm psf}_{a}$ (e.g., Paulin-Henriksson et al. 2008). Now we can write the two-point correlation of the shapes as,
@@
where we follow previous shape measurement literature in assuming $\boldsymbol{e}^{\rm true}$ does not correlate with $\boldsymbol{e}^{\rm psf}$, i.e., we assume $\langle\boldsymbol{e}^{\rm true}\boldsymbol{e}^{\rm psf}\rangle=0$. He
@@
for the same reasons mentioned above, $\boldsymbol{e}^{\rm true}$ does not correlate with $\boldsymbol{e}^{\rm psf}$ so the second and third term average to 0 given $\langle\boldsymbol{e}^{\rm true}\rangle=0$.
@@
where $\boldsymbol{e}^{\star}$ and $T^{\star}$ are the ellipticity and size of the light profile, respectively, of a star, while $\boldsymbol{e}^{\rm model}$ and $T^{\rm model}$ are the same for the PSF model evaluated at the location of the star. The coefficients $\alpha_{a},\beta_{a},\eta_{a}$ are measured using only the galaxies in tomographic bin $a$. The quantities $\boldsymbol{e}^{\rm model},\boldsymbol{e}^{\star}$, etc. are not dependent on tomographic bin, but the coefficients associated with a given term does depend on the bin. We can write the cross-term, $\langle\delta\boldsymbol{e}^{\rm psf}_{a}\delta\boldsymbol{e}^{\rm psf}_{b}\rangle$ as,
@@
We compute these bin-dependent coefficients using the same techniques as Paper I (see their Section 4.5) but now with only the source galaxies in tomographic bin $a$. The coefficient estimates are listed in Table 3, and are consistent with the non-tomographic case presented in Paper I. We estimate the PSF contamination to the shear two-point data vector, $\delta\xi_{\pm}^{\rm\,psf}$, using Equation (A.1) and the coefficients of Table 3. For each bin, we generate a distribution of $\delta\xi_{\pm}^{\rm\,psf}$ contributions. This is done by sampling from the joint posterior of the $\alpha,\beta,\eta$ parameters in that bin, and generate a $\delta\xi_{\pm}^{\rm\,psf}$ per sample. We then compute the $99.7\%$ upper/lower bounds on $\delta\xi_{\pm}^{\rm\,psf}$ and use those two estimates to contaminate a simulated data vector, where the latter is generated using the same approach as Section 5. The constraints from the contaminated data vectors are shown in Figure 7. The shifts in the cosmology are completely negligible, indicating that PSF contamination has negligible impact on the DECADE cosmology constraints.
@@ A.2 $B$-modes
In Paper I, we verify that the DECADE shape catalog does not contain any $B$-mode signal to within the statistical precision of the data. We now repeat this exercise, but perform the measurement for different tomographic bin combinations. Here, we have used the Namaster (Alonso et al. 2019) package to measure the $B$-modes power spectra in our data. Our measurement follows the same approach as Doux et al. (2022), and we also generate measurements covariances using the same, simulation-based approach of that work.
@@
Figure 8 shows a subset of our results. We find no evidence of $B$-modes in any tomographic bin combination. All measurements not shown in Figure 8 all have $p>0.3$. Given we have no detection of $B$-mode contamination, we do not estimate a $\delta\xi_{\pm}^{\,B{\rm-mode}}$ contribution and propagate the impact of a $B$-mode contaminant to cosmology. This follows the choices in Amon et al. (2022). Note that the harmonic-space estimator here differs from the real-space estimator used in Paper I. The former was chosen for this test as our original purpose was to contaminate the simulated data vector with $B$-modes (assuming a $B$-mode signal was detected) and this task requires as input the angular power spectra of $B$-modes. We have also verified that the non-tomographic measurement of the $B$-modes from Namaster is statistically consistent with no signal ($p=0.41$) which itself is consistent with the result from the real-space estimator in Paper I.
Appendix B Additional results from catalog splits
Run $a_{2}$ $\eta_{2}$ $\chi^{2}/N_{\rm dof}$ $p$
Airmass+ $1.2^{+2.8}_{-2.5}$ $2.1^{+1.9}_{-4.2}$ 233.1/220 0.260
Airmass- $2.0^{+4.7}_{-1.9}$ $3.23^{+0.81}_{-2.2}$ 268.1/220 0.015
Seeing+ $2.0^{+2.9}_{-2.1}$ $2.6^{+1.5}_{-3.5}$ 215.4/220 0.575
Seeing- $3.0^{+1.2}_{-1.0}$ $3.66^{+0.40}_{-0.64}$ 287.0/220 0.002
DCR-R.A.+ $2.7^{+1.8}_{-1.4}$ $3.25^{+0.78}_{-1.4}$ 234.6/220 0.238
DCR-R.A.- $0.6^{+1.8}_{-1.8}$ $1.6^{+2.5}_{-4.4}$ 252.5/220 0.066
DCR-Dec.+ $1.4^{+3.6}_{-2.2}$ $2.6^{+1.5}_{-4.0}$ 264.3/220 0.022
DCR-Dec.- $2.3^{+5.3}_{-1.9}$ $3.1^{+1.1}_{-2.1}$ 236.4/220 0.214
DCR+-$\rm e_{1}$ $2.1^{+2.6}_{-2.0}$ $2.9^{+1.1}_{-2.8}$ 245.9/220 0.111
DCR--$\rm e_{1}$ $2.2^{+1.6}_{-1.6}$ $3.06^{+0.98}_{-2.1}$ 233.2/220 0.259
DCR+-$\rm e_{2}$ $0.4^{+2.7}_{-2.3}$ $2.0^{+2.0}_{-4.4}$ 233.9/220 0.248
DCR--$\rm e_{2}$ $2.4^{+2.1}_{-1.7}$ $3.0^{+1.0}_{-2.4}$ 212.6/220 0.627
DECADE West $3.31^{+0.96}_{-0.72}$ $3.48^{+0.54}_{-0.81}$ 238.0/220 0.192
DECADE East $0.4^{+1.8}_{-1.6}$ $1.5^{+2.6}_{-5.2}$ 256.1/220 0.048
Pan-STARRS Calib. $2.9^{+1.1}_{-1.0}$ $3.52^{+0.50}_{-0.78}$ 256.8/220 0.045
SkyMapper Calib. $0.1^{+3.2}_{-3.1}$ $0.9^{+3.2}_{-4.4}$ 234.3/220 0.242
Skybright+ $1.4^{+2.6}_{-2.3}$ $2.1^{+1.9}_{-4.5}$ 262.1/220 0.027
Skybright- $1.2^{+2.0}_{-2.0}$ $1.9^{+2.2}_{-3.8}$ 251.7/220 0.070
Skysigma+ $2.5^{+6.5}_{-1.6}$ $3.19^{+0.84}_{-1.3}$ 265.3/220 0.020
Skysigma- $1.3^{+2.0}_{-2.2}$ $2.2^{+1.8}_{-4.4}$ 262.5/220 0.026
Exptime+ $1.9^{+1.9}_{-1.9}$ $3.2^{+1.0}_{-2.2}$ 268.0/220 0.015
Exptime- $1.7^{+1.9}_{-2.1}$ $1.5^{+2.5}_{-3.3}$ 204.6/220 0.764
Maglim+ $2.2^{+5.3}_{-1.9}$ $3.39^{+0.75}_{-1.3}$ 269.4/220 0.013
Maglim- $2.7^{+1.7}_{-1.3}$ $1.6^{+2.3}_{-2.7}$ 220.1/220 0.486
Nexp+ $2.0^{+5.2}_{-1.9}$ $3.45^{+0.68}_{-1.1}$ 262.6/220 0.026
Nexp- $2.4^{+1.9}_{-1.7}$ $1.6^{+2.5}_{-3.2}$ 221.3/220 0.462
Extinction+ $1.1^{+3.0}_{-2.2}$ $2.2^{+1.9}_{-3.6}$ 240.9/220 0.160
Extinction- $1.8^{+1.8}_{-1.8}$ $2.4^{+1.7}_{-3.2}$ 252.2/220 0.067
StarDensity+ $1.2^{+3.4}_{-2.2}$ $2.5^{+1.5}_{-3.9}$ 230.4/220 0.302
StarDensity- $2.4^{+1.7}_{-1.5}$ $2.8^{+1.2}_{-2.5}$ 239.6/220 0.173
Table 4: Constraints on two parameters from the TATT extensions to the IA model, the $\chi^{2}$ of the maximum posterior model, and the resulting p-value. Results are shown for all splits on survey property maps. The analogous results for the object property splits are in Table 5. Note that the posteriors are fairly non-Gaussian, and to enhance the interpretation of consistency/inconsistency of the results we utilize the $2\sigma$ (not $1\sigma$) bounds of the posterior. The p-values are not corrected for the “look elsewhere” effect. See text for details.
Run $a_{2}$ $\eta_{2}$ $\chi^{2}/N_{\rm dof}$ $p$
rz<0.85r-z<0.85 $1.2^{+4.2}_{-2.7}$ $2.7^{+1.4}_{-3.5}$ 253.7/220 0.059
rz>0.85r-z>0.85 $1.8^{+5.2}_{-2.1}$ $2.0^{+2.1}_{-4.8}$ 253.9/220 0.058
$m_{r}<22.5$ $1.3^{+5.2}_{-2.7}$ $2.7^{+1.4}_{-2.6}$ 271.9/220 0.010
$m_{r}>22.5$ $2.3^{+4.9}_{-1.8}$ $2.4^{+1.7}_{-2.7}$ 210.8/220 0.660
$m_{i}<22$ $0.9^{+3.2}_{-2.5}$ $2.1^{+2.0}_{-3.8}$ 215.1/220 0.581
$m_{i}>22$ $8^{+1.4}_{-1.2} 2.2$$2^{+1.8}_{-2.1}$ 217.5/220 0.534
$m_{z}<21.6$ $0.3^{+2.5}_{-2.3}$ $1.7^{+2.4}_{-4.5}$ 217.2/220 0.540
$m_{z}>21.6$ $2.7^{+5.8}_{-1.4}$ $3.11^{+0.93}_{-1.6}$ 196.6/220 0.869
SNR <27<27 $1.8^{+4.6}_{-2.3}$ $1.6^{+2.5}_{-3.9}$ 195.1/220 0.886
SNR >27>27 $0.9^{+1.7}_{-1.8}$ $1.9^{+2.1}_{-3.8}$ 228.4/220 0.335
T <0.48<0.48 $0.6^{+2.9}_{-2.8}$ $1.9^{+2.2}_{-4.7}$ 202.2/220 0.800
T >0.48>0.48 $2.8^{+1.4}_{-1.2}$ $3.0^{+1.0}_{-1.8}$ 249.7/220 0.082
Tpsf <0.45<0.45 $2.4^{+6.2}_{-1.7}$ $2.7^{+1.4}_{-2.2}$ 235.5/220 0.226
Tpsf >0.45>0.45 $2.9^{+1.3}_{-1.1}$ $3.65^{+0.42}_{-0.69}$ 256.1/220 0.048
T/Tpsf <1<1 $1.1^{+4.1}_{-2.9}$ $2.8^{+1.3}_{-3.8}$ 242.9/220 0.139
T/Tpsf >1>1 $3.10^{+1.2}_{-0.93}$ $2.9^{+1.1}_{-1.6}$ 242.0/220 0.148
Table 5: Similar to Table 4 but for splits performed using the object (not survey) properties. The p-values are not corrected for the “look elsewhere” effect. See text for details.
Figure 10: The mean redshift distributions, $n(z)$, for the four tomographic bins (columns) for some of the splits shown in Section 6.2 and 6.3. The fiducial redshift distribution is shown in gray. The text in the top shows the mean redshift, as well as the uncertainty on it as calibrated by the pipeline in Paper II. The text also denotes the split corresponding to each color. The statistics for the fiducial distributions are shown in top row. The area-based splits only lead to minimal changes in the mean $n(z)$ — though they may have greater effects in causing spatially varying $n(z)$, which will not be highlighted in this figure which presents the area-averaged $n(z)$ — while the object-based splits cause a clear separation in redshift. The mean redshifts for the fiducial case are,